OpenStudy (anonymous):

express as a sum of logarithms log v4 (38*19)

OpenStudy (lgbasallote):

hint \[\huge \log_2 (3 \times 4) \implies \log_2 3 + \log_2 4\] does that help?

OpenStudy (anonymous):

sadly no. i dont understand log that well in general and have no idea how to plug it into a calculator

OpenStudy (anonymous):

i have a ti-83 plus that i got from a friend, but im not too familiar with graphing calculators

OpenStudy (lgbasallote):

you don't need to use calculators here. just laws of logarithms

OpenStudy (lgbasallote):

\[\huge \log_a (bc) = \log_a b + \log_a c\] therefore... \[\huge \log_ 2 (3 \times 4) = \log_2 3 + \log_2 4\] making sense now?

OpenStudy (lgbasallote):

\[\huge \log_3 ( 2\times 4) = \log_3 2 + \log_ 3 4\]

OpenStudy (lgbasallote):

so now.. can you find \[\huge \log_4 (38 \times 19)\]

OpenStudy (anonymous):

yeah so i ended up with \[\log_{4}38+\log_{4} 19\] what next?

OpenStudy (lgbasallote):

that's it. that's the answer

OpenStudy (anonymous):

ok. thanks once again.

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

what about \[\ln \frac{ 8 }{ 9x ^{9}y}\]

OpenStudy (anonymous):

does it work the same way?

OpenStudy (lgbasallote):

what's the instruction?

OpenStudy (anonymous):

express in terms of sums and differences of logarithms

OpenStudy (lgbasallote):

sum works same way...but difference no \[\huge \ln (\frac ab) \implies \ln a - \ln b\] NOTE: \[\LARGE \ln (\frac a{bc}) \implies \ln a - (\ln b + \ln c) \implies \ln a - \ln b - \ln c\] do you get what i mean?

OpenStudy (anonymous):

k. i think i get it. thanks again

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

ugh.....everytime i think it gets easier it just gets harder.....can you help me with this too. I feel like i should be paying you for this. Express as a sum or difference of logarithms without exponents \[\log _{c5}\sqrt{\frac{ x ^{2} }{ y ^{5}z ^{7}}}\]

OpenStudy (lgbasallote):

without exponents...so this time you're going to use power rule...

OpenStudy (lgbasallote):

i think i should just give a similar example..it's hard to explain in words

OpenStudy (lgbasallote):

\[\large \ln \sqrt{\frac{a^3}{b^2c}} \implies \ln (\frac{a^3}{b^2 c})^{\frac 12} \implies \frac 12 \ln (\frac{a^3}{b^2 c}) \implies \frac 12 \ln a^3 - \frac 12 \ln b^2 - \frac 12 \ln c\] \[\large \implies \frac 32 \ln a- \frac 22 \ln b - \frac 12 ln c \implies \frac 32 \ln a - \ln b - \frac 12 \ln c\]

OpenStudy (lgbasallote):

the power rule is this: \[\huge \log_a b^c \implies c\log_a b\]